OpenStudy (anonymous):

determine the values of positive k for which the series from n=2 to infinity of ((n)^(1/2)) / (n-n^k)

OpenStudy (sirm3d):

is covergent? \[\large \frac{\sqrt n}{n-n^k}\frac{\sqrt n}{\sqrt n(\sqrt n-n^{k-1/2})}=\frac{1}{n^{1/2}-n^{k-1/2}}\] the series \[\large \sum \frac{1}{n^{1/2}-n^{n^k-1/2}}\] can be compared to a p-series that is convergent when \(p>1\). Here, \[\large p=k-1/2\]

OpenStudy (anonymous):

thank you so much. i was just really stuck on how to approach this. i'm confused on how you got to series 1/ (n^(1/2)- n^(n^k-1/2). i'm not sure where the from

OpenStudy (anonymous):

*where the n came from

OpenStudy (sirm3d):

factor n^(1/2), so that you can cancel it with the numerator.

OpenStudy (anonymous):

i saw that. it's just for your final summation you put an n to the n power. i think that was just a little mistake.

OpenStudy (anonymous):

how would you do a p series test if there are two ps under 1 fraction. you have n^(1/2) and n^(k-1/2)?